OpenStudy (babyslapmafro):

How do you come to the conclusion that sin^2xcscx=sinx ?

OpenStudy (anonymous):

since\[ \csc(x) = \frac{ 1 }{ \sin(x) }\] Then \[\sin^2(x)\times\frac{ 1 }{ \sin(x) } = \sin(x)\]

OpenStudy (babyslapmafro):

oh ok thanks, i was making it much more difficult than it needed to be...

OpenStudy (anonymous):

:)